OpenStudy (anonymous):

Short Answer: Simplify the following expression, and rewrite it in an equivalent form with positive exponents. (6x^-1)^-2

OpenStudy (anonymous):

6x^2

OpenStudy (anonymous):

(6x^-1)^-2 = (1/6x)^-2 = (6x)^2 = 36x^2